OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)?

OpenStudy (anonymous):

y=mx+b m=(10-4)/(4-2) = 6/2 = 3 solve for b

OpenStudy (anonymous):

y-y1 =m(x-x1 ) ----> we will use the point (1,2) = (x1,y1) here y-2=3(x-1)=3x-3 add 2 to both sides and you get your equation... y = 3x-1

OpenStudy (anonymous):

y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (anonymous):

y=3x-2

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

almost! instead of 2 its 1 y=3x-1 :)

OpenStudy (anonymous):

i mean not 1 y=3x-1......y=3x-1

OpenStudy (anonymous):